# Cost-aware Portfolios in a Large Universe of Assets

### v1

@@ 3 Simulation
First, the factor loadings $\bm{b}_{i}$ are drawn from a trivariate normal distribution $\mathsf{N}\left(\bm{\mu}_{b},\mathbf{cov}_{b}\right)$ with parameters specified in Table 1 below. Once generated, these factor loadings remain fixed throughout the simulations. The returns of the three factors over $n$ periods are also sampled from a trivariate normal distribution $\mathsf{N}\left(\bm{\mu}_{f},\mathbf{cov}_{f}\right)$, with the relevant parameters provided in Table 1. Throughout this simulation study, we assume that $\mathbb{E}(\mathbf{\bm{\epsilon}}\mid\mathbf{f})=\mathbf{0}$ and $\mathsf{Cov}(\bm{\epsilon}\mid\mathbf{f})=\operatorname{diag}\left(\sigma_{1}^{2},\cdots,\sigma_{p}^{2}\right)$. The standard deviations of the idiosyncratic errors $\varepsilon_{i}$ are generated from a gamma distribution with a shape parameter of 3.3586 and a scale parameter of 0.1876. Then, the generated parameters $\left\{\sigma_{i}\right\}$ remain fixed across simulations. Finally, the idiosyncratic errors are generated from a normal distribution, where the standard deviations are set to $\left\{\sigma_{i}\right\}$.
@@
To examine the performance of the proposed method, we include the following three models for comparison in the simulation study.
  • 1.
    The conventional Markowitz mean-variance model
    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1},\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.1)
    where the $w_{1,i}$ is denoted as the $i$th element of $\bm{w}_{1}$.
  • 2.
    The regularized mean-variance model with $\ell_{1}$ penalty.
    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{t}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\lambda\|\bm{w}_{1}\|_{1},\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.2)
  • 3.
    The mean-variance model with the transaction cost.
    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}_{1}),\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.3)
The models described above are for comparison during the first portfolio construction stage. By replacing $\bm{w}$ with $\bm{\delta}$ and substituting the subscript $1$ with $t$, the corresponding three models for the reallocation stage can be obtained.
We set $p=2000$, $m=5$ and $n=200$, dividing the data into five stages to calculate the out-of-sample annual Return, Cost, Turnover, Leverage and Sharpe ratio. The optimal tuning parameter $\lambda^{opt}$ is selected based on the highest in-sample Sharpe ratio.
The main method is the proposed Cost-Aware Portfolio Estimator (CAPE). For comparison, we utilize three different baseline models: the Mean-Variance (MV) model (3.1), the Penalized Mean-Variance model (PMV) (3.2), and the Cost-Aware Mean-Variance (CMV) model (3.3) to compute portfolio weights at each stage. We employ the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE) used in Ledoit and Wolf (2017) to estimate the covariance matrix $\bm{\Sigma}$. Specifically, we utilize the functions nonlin_shrinkLW() and cov.shrink() from the R packages HDShOP and corpcor to calculate NLSE and LSE, respectively. The results from the NLSE will be presented in the main text, while the results from the LSE are included in the Supplementary Materials.
@@
S2 MV 8.158(0.273) 3.370(0.004) 16.494(0.010) 7.747(0.005) 1.004(0.034)
PMV 7.911(0.139) 0.740(0.002) 4.363(0.006) 1.681(0.003) 2.020(0.036)
CMV 8.330(0.244) 2.693(0.003) 14.740(0.008) 6.870(0.004) 1.162(0.034)
CAPE 7.881(0.128) 0.616(0.002) 3.855(0.005) 1.428(0.003) 2.196(0.036)
S3 MV 4.138(0.273) 6.482(0.024) 22.684(0.040) 7.754(0.005) 0.436(0.029)
PMV 8.107(0.162) 1.056(0.004) 5.275(0.009) 2.799(0.003) 1.747(0.035)
CMV 9.634(0.246) 4.683(0.015) 19.265(0.029) 6.667(0.004) 1.201(0.031)
CAPE 8.189(0.152) 0.839(0.003) 4.516(0.008) 2.458(0.003) 1.908(0.035)
S4 MV 4.285(0.275) 6.541(0.024) 22.791(0.040) 7.736(0.005) 0.451(0.029)
PMV 8.472(0.173) 1.193(0.004) 5.709(0.011) 3.395(0.003) 1.728(0.036)
CMV 11.748(0.247) 4.564(0.014) 19.036(0.027) 6.527(0.004) 1.476(0.031)
CAPE 8.554(0.162) 0.950(0.004) 4.882(0.009) 3.034(0.003) 1.874(0.036)
S5 MV 4.397(0.271) 6.517(0.024) 22.747(0.041) 7.742(0.005) 0.466(0.029)
PMV 9.036(0.173) 1.244(0.005) 5.902(0.012) 3.733(0.004) 1.804(0.035)
CMV 12.730(0.245) 4.477(0.013) 18.860(0.026) 6.476(0.004) 1.609(0.031)
CAPE 9.137(0.163) 0.994(0.004) 5.057(0.010) 3.369(0.004) 1.954(0.035)
  • Notes: MV, PMV, CMV and CAPE are mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with nonlinear shrinkage estimated covariance matrices), respectively. S1 is the portfolio construction stage, S2-S5 are rebalancing stages, respectively. And the values in columns of Return, Cost, Turnover, and Leverage are in percentages. SR is the Sharpe ratio post transaction costs.
@@
S2 MV 9.879(0.273) 1.649(0.001) 16.494(0.010) 7.747(0.005) 1.284(0.036)
PMV 8.214(0.139) 0.436(0.001) 4.363(0.006) 1.681(0.003) 2.117(0.036)
CMV 9.801(0.268) 1.578(0.001) 15.783(0.010) 7.391(0.005) 1.304(0.036)
CAPE 8.132(0.131) 0.383(0.001) 3.828(0.005) 1.414(0.003) 2.224(0.036)
S3 MV 8.352(0.271) 2.268(0.004) 22.684(0.040) 7.754(0.005) 1.063(0.035)
PMV 8.635(0.162) 0.528(0.001) 5.275(0.009) 2.799(0.003) 1.889(0.036)
CMV 8.448(0.267) 2.171(0.004) 21.713(0.038) 7.610(0.005) 1.092(0.034)
CAPE 8.572(0.155) 0.455(0.001) 4.551(0.008) 2.449(0.003) 1.970(0.036)
S4 MV 8.547(0.275) 2.279(0.004) 22.791(0.040) 7.736(0.005) 1.092(0.035)
PMV 9.094(0.173) 0.571(0.001) 5.709(0.011) 3.395(0.003) 1.890(0.036)
CMV 8.660(0.270) 2.198(0.004) 21.979(0.039) 7.612(0.005) 1.121(0.035)
CAPE 8.994(0.166) 0.494(0.001) 4.945(0.010) 3.037(0.003) 1.947(0.036)
S5 MV 8.639(0.271) 2.275(0.004) 22.747(0.041) 7.742(0.005) 1.105(0.035)
PMV 9.690(0.173) 0.590(0.001) 5.902(0.012) 3.733(0.004) 1.971(0.035)
CMV 8.708(0.270) 2.197(0.004) 21.974(0.040) 7.627(0.005) 1.128(0.035)
CAPE 9.610(0.167) 0.514(0.001) 5.135(0.011) 3.386(0.004) 2.030(0.036)
@@
To emulate real-world investment scenarios, we report the performance of portfolios across four stages (S2-S5). The first stage involves forming the initial portfolio, akin to establishing a new fund at its inception. The second stage corresponds to the first rebalancing, while the third stage entails the second rebalancing, and so forth. Following the conventions established in the literature, we define the parameters for proportional and quadratic transaction costs as $\bm{\beta}=\beta\times\mathbf{1}_{p}$ and $\bm{\alpha}=\alpha\times\mathbf{1}_{p}$, where $\beta=0.15$ and $\alpha=0.001$ respectively. The costs in the following tables are therefore interpreted as the percentage of the normalized initial wealth that needs to cover the costs of transactions. Additionally, we quantify portfolio turnover and leverage at each stage $k=2,3,4,5$. The definitions of these metrics are as follows:
$\text{Turnover}_{k}=\sum\limits_{i=1}^{p}|w_{k,i}-w^{+}_{k-1,i}|,w_{1^{+},i}=0,$
@@
The simulation results are summarized in Tables 2 and 3. The classic Markowitz mean-variance portfolio exhibits the highest turnover, resulting in the largest transaction costs and the lowest out-of-sample Sharpe ratio. In contrast, for both the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE), our proposed method, CAPE, demonstrates the highest Sharpe ratio and the lowest transaction costs. This finding underscores the importance of incorporating costs into the optimization problem.
@@ 4.1 S&P 500 Index
We collected daily S&P 500 Index composition stocks data covering the holding period from 02-Jan-2017 to 23-Dec-2020. Our portfolio is rebalanced every 251 trading days, resulting in three decision days and four stages S1-S4 (construction, and then three rebalances). The out-of-sample period extends from 01-Jan-2018 to 23-Dec-2020. On each decision day, the covariance matrix $\bm{\Sigma}$ is estimated using two different methods: the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE), both based on the most recent 251 daily returns. In the main text we show the results using NLSE. The similar results obtained using the LSE are presented in Section S3 of the supplementary materials.
In the study we consider those stocks with complete records over the holding period. Specifically, on the first decision day, a stock has a complete set of 251 days of historical returns, and by the last decision day, it also possesses a full set of returns for the subsequent 251 days. As a result, our stock pool comprises 462 stocks.
In addition to our proposed CAPE method and the three comparison methods (MV, PMV and CMV) shown in the simulation studies, we also include the benchmark equally-weighted portfolio (DeMiguel et al., 2009), where the weights are given as $1/p$. Following the convention of the literature we just label it as ‘1/N’ portfolio in the empirical tables. We assess several performance metrics, including out-of-sample annual return, transaction costs, turnover, leverage, out-of-sample risk, the annualized Sharpe ratio (we report it separately for stages S2-S4) for holding the portfolio at its current rebalanced composition until the next rebalancing, and the overall Sharpe ratio during the whole S2-S4 period. The performances of the various methods are reported in Tables 4 and 5, which present results for proportional and quadratic transaction costs, respectively.
Our proposed CAPE method demonstrates the best overall Sharpe ratio performance during the S2 to S4 phases. In terms of annual returns, the CAPE method has shown the ability to control losses during the period (S2) of overall declines in the S&P 500 and maintaining substantial gains at the lower level of risk in the period of overall profitability (S3-S4). Excluding the equally-weighted (‘1/N’) portfolio, the CAPE method exhibits one of the lowest turnover among actively managed portfolios. More importantly, CAPE strikes a balance between the required return and risk, considering transaction costs. The CAPE is usually the best performer in annual Sharpe ratio, and has the best Sharpe ratio in overall out-of-sample period S2-S4. Furthermore, the CAPE method is more robust than the other comparison methods concerning different transaction cost functions.
Table 4: S&P 500 quadratic transaction costs, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
@@
1/N S2 -3.835 0.033 1.000 0.000 -0.244 0.863
S3 29.247 0.002 0.181 0.000 2.336
S4 47.154 0.002 0.153 0.000 1.067
MV S2 -2.797 1.034 4.382 1.691 -0.178 0.883
S3 29.977 2.474 6.754 2.148 2.404
S4 44.761 1.878 5.878 1.755 1.052
PMV S2 -6.943 0.826 3.550 1.275 -0.466 0.704
S3 24.586 0.796 1.859 1.530 2.387
S4 37.516 0.291 0.780 1.460 0.905
CMV S2 -2.189 0.808 3.874 1.437 -0.145 0.755
S3 42.578 3.109 7.611 2.799 2.614
S4 46.451 2.603 6.920 2.385 0.742
CAPE S2 -2.016 0.818 3.444 1.222 -0.133 0.970
S3 30.755 1.183 3.226 1.626 2.822
S4 57.831 0.588 1.842 1.433 1.204
  • Notes: ‘1/N’, MV, PMV, CMV and CAPE are the equally-weighted, mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with nonlinear shrinkage estimated covariance matrices), respectively. S1 is the portfolio construction stage, S2-S4 are rebalancing stages, respectively. And the values in columns of Return and Cost represent percentages. The “overall SR” in the last column is the post-transaction cost Sharpe ratio of the portfolio for holding it from S2-S4 to show the overall performance in our sample period. More details of other performance measure can be found in the notes of Table 2.
Table 5: S&P 500 proportional transaction costs, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.

### v2

@@ 3 Simulation
First, the factor loadings $\mathbf{b}_{i}$ are drawn from a trivariate normal distribution ${N}\left({\bm{\mu}}_{b},\mathbf{cov}_{b}\right)$ with parameters specified in Table 1 below. Once generated, these factor loadings remain fixed throughout the simulations. The returns of the three factors over $n$ periods are also sampled from a trivariate normal distribution ${N}\left({\bm{\mu}}_{f},\mathbf{cov}_{f}\right)$, with the relevant parameters provided in Table 1. Throughout this simulation study, we assume that $\mathbf{E}(\bm{\varepsilon}\mid\mathbf{f})=\mathbf{0}$ and $\mathbf{Cov}(\bm{\varepsilon}\mid\mathbf{f})$ is a diagonal matrix with entries $\sigma_{1}^{2},\dots,\sigma_{p}^{2}$ on the diagonal. The standard deviations of the idiosyncratic errors $\varepsilon_{i}$ are generated from a gamma distribution with a shape parameter of 3.3586 and a scale parameter of 0.1876. Then, the generated parameters $\left\{\sigma_{i}\right\}$ remain fixed across simulations. Finally, the idiosyncratic errors are generated from a normal distribution, where the standard deviations are set to $\left\{\sigma_{i}\right\}$.
@@
We set $p=2000$, $m=5$ and $n=200$, dividing the data into five stages to calculate the out-of-sample annual Return, Cost, Turnover, Leverage and Sharpe ratio. And S0 is the portfolio construction stage, S1-S4 are rebalancing stages, respectively. The optimal tuning parameter $\lambda^{opt}$ is selected based on the highest in-sample Sharpe ratio.
The main method is the proposed Cost-Aware Portfolio Estimator under SCAD Penalty (CAPE-S). For comparison, we utilize three different baseline models montioned in Section 2.1: the classical Mean-Variance (MV) model (2.1), the Penalized Mean-Variance (PMV) model (2.2), and the Cost-Aware Mean-Variance (CMV) model (2.3) to compute portfolio weights at each stage. We estimate the covariance matrix $\bm{\Sigma}_{t}$ using the linear shrinkage estimator (LSE; Ledoit and Wolf (2004)) and the nonlinear shrinkage estimator (NLSE; Ledoit and Wolf (2020)), while the mean vector $\bm{\mu}_{t}$ is estimated by the sample mean. The results based on the LSE, which satisfy Assumptions (A2) and (A3) outlined earlier, are presented in the main text. In contrast, the results based on the NLSE, which satisfy only Assumption (A3) but not (A2), are reported in the Supplementary Materials. Despite not meeting (A2), the NLSE performs well in both simulation studies and real data applications.
@@
S1 MV 7.992(0.241) 4.120(0.005) 18.160(0.010) 8.580(0.005) 1.044(0.032)
PMV 7.766(0.107) 0.624(0.001) 4.082(0.006) 1.541(0.003) 2.599(0.036)
CMV 8.285(0.212) 2.993(0.003) 15.693(0.009) 7.347(0.004) 1.279(0.033)
CAPE-S 7.723(0.097) 0.500(0.001) 3.562(0.005) 1.281(0.002) 2.851(0.036)
S2 MV 3.581(0.242) 7.736(0.024) 18.161(0.011) 8.580(0.005) 0.374(0.025)
PMV 8.269(0.125) 0.831(0.002) 4.845(0.008) 2.662(0.003) 2.290(0.035)
CMV 9.635(0.215) 4.974(0.013) 20.216(0.026) 7.186(0.004) 1.294(0.029)
CAPE-S 8.247(0.115) 0.631(0.002) 4.088(0.007) 2.291(0.003) 2.506(0.035)
S3 MV 3.670(0.250) 7.792(0.024) 18.131(0.010) 8.566(0.005) 0.385(0.026)
PMV 8.785(0.139) 0.934(0.003) 5.234(0.009) 3.305(0.003) 2.269(0.036)
CMV 11.613(0.220) 4.870(0.012) 20.012(0.025) 7.050(0.004) 1.572(0.030)
CAPE-S 8.758(0.128) 0.708(0.002) 4.406(0.007) 2.894(0.003) 2.469(0.036)
S4 MV 3.764(0.245) 7.787(0.025) 18.148(0.010) 8.574(0.005) 0.396(0.026)
PMV 9.312(0.137) 0.981(0.003) 5.419(0.010) 3.692(0.004) 2.332(0.035)
CMV 12.530(0.215) 4.797(0.012) 19.855(0.024) 6.997(0.004) 1.705(0.030)
CAPE-S 9.294(0.128) 0.745(0.002) 4.566(0.008) 3.269(0.003) 2.530(0.035)
  • Notes: MV, PMV, CMV and CAPE-S are mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with linear shrinkage estimated covariance matrices), respectively. S0 is the portfolio construction stage, S1-S4 are rebalancing stages, respectively. And the values in columns of Return and Cost represent percentages.
@@
S1 MV 10.296(0.241) 1.816(0.001) 18.160(0.010) 8.580(0.005) 1.489(0.035)
PMV 7.983(0.107) 0.408(0.001) 4.082(0.006) 1.541(0.003) 2.698(0.036)
CMV 10.169(0.235) 1.729(0.001) 17.293(0.010) 8.147(0.005) 1.513(0.035)
CAPE-S 7.879(0.099) 0.353(0.001) 3.527(0.005) 1.264(0.003) 2.878(0.036)
S2 MV 8.833(0.241) 2.484(0.004) 24.840(0.039) 8.580(0.005) 1.245(0.034)
PMV 8.615(0.125) 0.484(0.001) 4.845(0.008) 2.662(0.003) 2.420(0.035)
CMV 8.965(0.237) 2.368(0.004) 23.684(0.037) 8.408(0.005) 1.287(0.034)
CAPE-S 8.471(0.117) 0.410(0.001) 4.097(0.007) 2.277(0.003) 2.552(0.036)
S3 MV 8.968(0.248) 2.494(0.004) 24.937(0.038) 8.566(0.005) 1.267(0.035)
PMV 9.196(0.139) 0.523(0.001) 5.233(0.009) 3.305(0.003) 2.414(0.037)
CMV 9.078(0.245) 2.397(0.004) 23.967(0.038) 8.415(0.005) 1.304(0.035)
CAPE-S 9.023(0.131) 0.444(0.001) 4.438(0.008) 2.892(0.003) 2.523(0.037)
S4 MV 9.060(0.244) 2.492(0.004) 24.918(0.040) 8.574(0.005) 1.278(0.035)
PMV 9.751(0.137) 0.542(0.001) 5.419(0.010) 3.692(0.004) 2.483(0.035)
CMV 9.142(0.242) 2.399(0.004) 23.986(0.040) 8.432(0.005) 1.310(0.035)
CAPE-S 9.586(0.130) 0.461(0.001) 4.613(0.008) 3.282(0.003) 2.586(0.036)
@@
To emulate real-world investment scenarios, we report the performance of portfolios across four stages (S1-S4). The first stage involves forming the initial portfolio, akin to establishing a new fund at its inception. The second stage corresponds to the first rebalancing, while the third stage entails the second rebalancing, and so forth. Following the conventions established in the literature, we define the parameters for proportional and quadratic transaction costs as $\bm{\beta}=\beta\times\mathbf{1}_{p}$ and $\bm{\alpha}=\alpha\times\mathbf{1}_{p}$, where $\beta=0.15$ and $\alpha=0.001$ respectively. The costs in the following tables are therefore interpreted as the percentage of the normalized initial wealth that needs to cover the costs of transactions. Additionally, we quantify portfolio turnover and leverage at each stage $k=1,2,3,4$. The definitions of these metrics are as follows:
$\text{Turnover}_{k}=\sum\limits_{i=1}^{p}|w_{k,i}-w^{+}_{k-1,i}|\quad\text{and}\quad w^{+}_{0,i}=0,$
@@
We evaluate the performance of CAPE-S and three benchmark models under both quadratic and proportional transaction cost settings. The results, reported in Tables 2 and 3, demonstrate the consistent superiority of our proposed method in balancing return, risk, and cost-efficiency across all rebalancing stages (S1–S4).
Under quadratic transaction costs (Table 2), CAPE-S yields the highest Sharpe ratio at every stage, significantly outperforming the traditional Mean-Variance (MV), Penalized Mean-Variance (PMV), and Cost-aware Mean-Variance (CMV) models. For example, in stage S1, CAPE-S achieves a Sharpe ratio of 2.851, compared to 2.599 (PMV), 1.279 (CMV), and 1.044 (MV). Notably, CAPE-S accomplishes this with the lowest transaction cost (0.500%) and lowest turnover (3.562%), indicating efficient reallocation with minimal trading. This pattern persists through stages S2 to S4, with CAPE-S consistently attaining higher return-to-risk ratios, while incurring substantially less cost and leverage than the MV and CMV models.
Similar conclusions hold under proportional transaction costs (Table 3). CAPE-S again delivers the best Sharpe ratio in each stage, with performance gains most prominent in S1 and S4, where the Sharpe ratios reach 2.878 and 2.586, respectively. Although PMV also exhibits cost-awareness, it sacrifices return and remains less effective in capturing the full benefit of transaction cost control. Compared to CMV and MV, which often yield higher raw returns, CAPE-S achieves better overall risk-adjusted performance by avoiding excessive leverage and unnecessary trading. For instance, in S2, CAPE-S reduces the turnover to 4.097% and the cost to 0.410%, about 16.5% of the MV model’s transaction cost, while increasing the Sharpe ratio by approximately 105%.
These findings highlight the advantage of our model in adapting to real-world trading frictions, demonstrating its ability to attain high returns while tightly controlling costs and maintaining portfolio sparsity. CAPE-S successfully balances the trade-off between return and cost, validating its practical value for large-scale portfolio management.
@@ 4 Real Data Application
In this section, we assess the empirical performance of the proposed CAPE-S method using two major U.S. equity universes: the S&P 500 and the Russell 2000 indices. For each index, we collect daily return data over the investment horizon from 02-Jan-2017 to 23-Dec-2020, and implement portfolio rebalancing every 251 trading days. This results in three decision points and four stages: an initial formation period (S0), followed by three rebalancing stages (S1–S3). The out-of-sample evaluation period spans from 01-Jan-2018 to 23-Dec-2020.
A key feature of this empirical exercise is to use real transaction cost data instead of the pre-determined constant costs coefficients. First, to ensure consistency, we cleaned the data by removing stocks with incomplete return histories or missing auxiliary data required for real transaction cost estimation (e.g., bid-ask spreads, trading volumes). As a result, we retain 457 stocks from the S&P 500 and 935 stocks from the Russell 2000 for analysis. The sample of stocks remain broadly representative of their respective index compositions.
At each decision point, the covariance matrix $\bm{\Sigma}_{t}$ is estimated using two methods: the linear shrinkage estimator (LSE) and the nonlinear shrinkage estimator (NLSE), both based on the most recent 251 trading days. Similar to the setting described in Section 3, we report the results based on the LSE in the main text, while the results obtained from the NLSE are provided in Section 3 of the Supplementary Materials.
Our evaluation includes the CAPE-S method, three benchmark approaches—mean-variance (MV), penalized mean-variance (PMV), and cost-aware mean-variance (CMV)—and the equally-weighted “1/N” portfolio (DeMiguel et al., 2009). We compare their performance using several metrics: annual return, transaction cost, turnover, leverage, stage-specific Sharpe ratios (S1–S3), and the overall Sharpe ratio across the entire out-of-sample period.
In contrast to the simulation studies, where transaction cost coefficients were fixed across assets (i.e., 0.15 for quadratic costs and 0.001 for proportional costs), such fixed cost specification is likely unrealistic in real-world applications. In empirical settings, transaction costs often vary substantially across stocks due to liquidity, trading frequency, and other microstructure factors. Therefore, we adopt asset-specific transaction cost coefficients for both proportional and quadratic forms, as discussed in the methodology sections.
For proportional transaction cost coefficients, we follow the empirical estimates provided in Renò et al. (2025), who use the bid-ask spread estimator proposed by Hasbrouck (2009) to proxy the proportional transaction costs at the individual stock level. These estimates are averaged over the formation window and are available for a wide range of U.S. equities from 1927 to the present, including both S&P 500 and Russell 2000 constituents.
For quadratic transaction cost coefficients, we define the asset-specific cost coefficient as twice the square of the corresponding proportional cost estimate. This specification ensures consistency in scaling and penalizes large rebalancing moves more heavily for less liquid stocks. Moreover, it reflects the increasing marginal cost nature of market impact, thereby capturing a more realistic trading environment.
By incorporating heterogeneity in cost parameters, our empirical analysis more accurately reflects the practical implementation of portfolio strategies under realistic trading frictions.
@@ 4.1 S&P 500 Index
We first present the empirical results on the S&P 500 dataset. The performance of all methods under both quadratic and proportional transaction cost models is summarized in Tables 4 and 5, respectively.
Table 4: S&P 500 with quadratic transaction cost.
@@
1/N S1 -3.666 0.017 1.000 0.000 -0.234 0.869
S2 29.236 0.001 0.181 0.000 2.341
S3 47.369 0.002 0.154 0.000 1.073
MV S1 -7.313 2.545 9.794 4.397 -0.329 0.772
S2 36.261 13.795 19.462 7.166 1.254
S3 141.735 12.551 18.118 6.441 1.158
PMV S1 -10.570 1.346 6.082 2.541 -0.570 0.665
S2 23.493 1.569 3.081 2.867 2.150
S3 78.677 0.532 1.561 2.796 1.027
CMV S1 -5.150 1.488 7.536 3.268 -0.266 0.891
S2 48.870 8.571 15.235 6.338 2.081
S3 96.660 6.899 13.651 5.504 1.127
CAPE-S S1 -3.687 1.631 6.951 2.976 -0.188 0.915
S2 30.491 2.349 4.766 3.265 2.490
S3 77.174 0.778 2.283 2.873 1.257
  • Notes: ‘1/N’, MV, PMV, CMV and CAPE are the equally-weighted, mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with linear shrinkage estimated covariance matrices), respectively. S1 is the portfolio construction stage, S2–S3 are rebalancing stages, respectively. And the values in columns of Return and Cost represent percentages.
Table 5: S&P 500 with proportional transaction cost.
@@
Our proposed CAPE-S method achieves the highest overall out-of-sample Sharpe ratio across the three stages (S1 to S3) under both quadratic and proportional transaction cost settings. Examining performance at the stage level provides further insights:
  • Stage S1 (2018) coincided with a period of heightened market volatility and an overall downturn, particularly during Q4 2018 when the S&P 500 experienced a sharp correction. While all active strategies posted negative returns, CAPE-S recorded the smallest loss among them under the proportional cost setting (-2.694%) and one of the smallest losses under the quadratic cost setting (-3.687%), demonstrating resilience during stressed market conditions. Turnover (7.147% proportional, 6.951% quadratic) and leverage (3.074 proportional, 2.976 quadratic) remained moderate relative to other active strategies.
  • Stage S2 (2019) represented a strong recovery year, with broad-based gains across the S&P 500 driven by accommodative monetary policy and easing trade tensions. Under both cost structures, CAPE-S delivered competitive returns (37.775% proportional, 30.491% quadratic) and achieved one of the highest Sharpe ratios among actively managed portfolios (1.666 proportional, 2.490 quadratic), while maintaining transaction costs (4.128% proportional, 2.349% quadratic) and turnover (15.688% proportional, 4.766% quadratic) at moderate levels compared to MV, PMV, and CMV.
  • Stage S3 (2020) covered a highly turbulent market cycle, including the COVID-19-induced crash in Q1 2020 followed by a rapid rebound. CAPE-S generated solid returns (108.519% proportional, 77.174% quadratic) and achieved the highest Sharpe ratio under the proportional cost setting (1.263), while keeping leverage (6.007 proportional, 2.873 quadratic) relatively contained compared to MV and CMV. This highlights its adaptability to sharp market reversals.
Across all stages, excluding the passive 1/N portfolio, CAPE-S consistently maintains one of the lowest turnover levels among actively managed strategies, reflecting operational efficiency. Its leading overall Sharpe ratio under both cost specifications (0.915 quadratic, 0.909 proportional) highlights its capability to balance returns, risk, and transaction costs, demonstrating robustness to varying cost structures.